# Planar Juggling of a Devil-Stick using Discrete VHCs

### v1

@@ V-A Orbit Selection
A distinct stable juggling motion described by DVHCs (12) and velocity constraints (21), with $\theta^{\mathrm{odd}}$ and $\theta^{\mathrm{even}}$ chosen to satisfy (39), is the orbit [^2] The orbit can equivalently be specified based on $\omega_{k}$ $\forall\,\theta_{k}=\theta^{\mathrm{even}}$.
@@ V-B Orbital Stabilization
If $(q,\dot{q})\in\mathcal{O}^{*}$, the system trajectory lies on $\mathcal{O}^{*}$ under the inputs from (27) and (28) alone. Therefore, the intersection of $\mathcal{O}^{*}$ with $\Sigma$ is a fixed point $z(j)=z^{*}$, $I(j)=I^{*}$, $r(j)=r^{*}$ of the map $\mathbb{P}$; $I^{*}$ and $r^{*}$ are obtained from (45) and (46) respectively:
@@
If $(q,\dot{q})\notin\mathcal{O}^{*}$, the inputs (27) and (28) drive system trajectories to an orbit such that $\rho_{k}=0$, but not necessarily the desired orbit $\mathcal{O}^{*}$. We stabilize $\mathcal{O}^{*}$ by using the ICPM approach [16] to stabilize the fixed point $z^{*}$ on $\Sigma$. First, linearize $\mathbb{P}$ about $z(j)=z^{*}$ and $I(j)=I^{*},\,r(j)=r^{*}$ as
@@
where $\mathcal{A}\in\mathbb{R}^{5\times 5}$, $\mathcal{B}\in\mathbb{R}^{5\times 2}$ may be computed numerically [11]. If the pair $(\mathcal{A},\mathcal{B})$ is controllable, $\mathcal{O}^{*}$ is asymptotically stabilized by the discrete feedback
@@ VI-A Enforcing the DVHC
We consider the DVHC in (12) with
$\alpha=0.6131,\quad\beta=3$ (57)
@@
$\theta^{\mathrm{odd}}=\pi/6,\quad\theta^{\mathrm{even}}=\pi-\theta^{\mathrm{odd}}=5\pi/6$ (58)
which ensures that the DZD in (38) permits stable, 2-periodic orbits - see Theorem 1. The resulting value of ${\Delta\theta^{*}}=\pi/3$. The control inputs are obtained using (27) and (28) with
@@
which agree with (43) and (44). The values of $\omega_{k}$ and $\delta_{k}$ are plotted in Fig.2(e) and Fig.2(f). The control inputs $I_{k}$ and $r_{k}$ are shown in Fig.2(g) and Fig.2(h) respectively. For the 2-periodic juggling motion in (61), $I_{k}=-(-1)^{k}\times 0.5890$ Ns and $r_{k}=0.0308$ m, in accordance with (45) and (46). The trajectory of the center-of-mass of the devil-stick corresponding to these results is shown in Fig.3(a).
@@
Figure 3: Trajectory of the center-of-mass of the devil-stick corresponding to the results in (a) Fig. 2 showing 2-periodic juggling, and (b) Fig. 4 showing planar symmetric juggling.
VI-B Stabilization of a Periodic Orbit
@@
We now demonstrate the effectiveness of the controller described in Section V in stabilizing a desired orbit, specifically the orbit corresponding to planar symmetric juggling in [8, 9]. We consider the DVHC in (12) with the parameters in (57) and the same choices of $\theta^{\mathrm{odd}}$ and $\theta^{\mathrm{even}}$ in (58). Of the infinitely many stable, 2-periodic orbits, we choose to stabilize the orbit for which $\omega^{*}=-4.1888$, obtained from (47). This results in $\delta_{k}=0.5$ s $\forall k$ corresponding to stable juggling. This orbit can be described using (48) as
$\begin{split}\mathcal{O}^{*}=\{(q,\dot{q}):h(k)&=\Phi(\theta_{k}),v(k)=\Psi(\theta_{k},\omega_{k}),\\ \omega_{k}&=-4.1888\,\,\forall\,\theta_{k}=\pi/3\}\end{split}$ (62)
@@
$\Sigma=\{(q,\dot{q})\in\mathbb{R}^{2}\times S^{1}\times\mathbb{R}^{3}:\theta=\pi/3,\omega<0\}$ (63)
@@
The matrices $\mathcal{A}$ and $\mathcal{B}$ in (53) are
@@
The same initial conditions (60) are used, and the simulation results are shown in Fig.4 for $k=1$ through $k=20$, a duration of approximately $9.59$ s. The components of $\rho_{k}$ are shown in Fig.4(a)-(b), and the components of $D\rho_{k}$ are shown in Fig.4(c)-(d). It is seen that the system trajectory converges to $\rho_{k}=0$, and $D\rho_{k}\to 0$ as $\rho_{k}\to 0$. The values of $\omega_{k}$, shown in Fig.4(e), converge to values in agreement with (47), $\omega^{*}=-4.1888$ rad/s, indicating stabilization of the orbit in (62). The time-of-flight, plotted in Fig.4(f), converges to the constant value $\delta_{k}=0.5$ s. The control inputs $I_{k}$ and $r_{k}$ are shown in Fig.4(g) and Fig.4(h) respectively. The inputs are obtained from (56) for $j=1,2,\dots,7$, i.e., $k=1,3,\dots,13$. The controller $u(j)$ is inactive for $j>7$ as the system trajectory is sufficiently close to $\mathcal{O}^{*}$. For planar symmetric juggling, $I_{k}=-(-1)^{k}\times 0.5664$ Ns and $r_{k}=0.0308$ m in accordance with (45) and (46). The trajectory of the center-of-mass of the devil-stick corresponding to these results is shown in Fig.3(b).
VII Conclusion
The problem of juggling a devil-stick alternately between two angular configurations was addressed using DVHCs. This paper extends the work in [11], which considered rotations of the passive coordinate to realize propeller motion, to oscillations of the passive coordinate to realize planar juggling. The DVHC approach results in a rich set of juggling motions with states that are not necessarily symmetric about the vertical axis; the results for planar symmetric juggling in [8] are shown to be a special case. It is demonstrated that DVHCs are an effective tool for design and control of the trajectory of an object juggled using impacts. Similar to the zero dynamics induced by VHCs, the DZD can be used to infer the nonlinear stability characteristics of the hybrid system subject to DVHCs, without relying on linearization about the fixed point of a Poincaré map [9]. Future work will focus on exploring a larger class of DVHCs for juggling systems and extension to other underactuated systems with impulsive inputs.

### v2

@@ V-A Orbit Selection
A distinct stable juggling motion described by DVHCs satisfying (39) is the orbit [^2] The orbit can equivalently be specified based on $\omega_{k}$ $\forall\,\theta_{k}=\theta^{\mathrm{even}}$.
@@ V-B Orbital Stabilization
We recall the approach to orbit stabilization in [khandelwal_discrete_2025]. If $(q,\dot{q})\in\mathcal{O}^{*}$, the system trajectory lies on $\mathcal{O}^{*}$ under the inputs from (27) and (28) alone. Therefore, the intersection of $\mathcal{O}^{*}$ with $\Sigma$ is a fixed point $z(j)=z^{*}$, $I(j)=I^{*}$, $r(j)=r^{*}$ of the map $\mathbb{P}$; $I^{*}$ and $r^{*}$ are obtained from (47) and (48):
@@
If $(q,\dot{q})\notin\mathcal{O}^{*}$, the inputs (27) and (28) drive system trajectories to an orbit such that $\rho_{k}=0$, but not necessarily the desired orbit $\mathcal{O}^{*}$. We stabilize $\mathcal{O}^{*}$ by using the ICPM approach [kant_orbital_2020] to stabilize the fixed point $z^{*}$ on $\Sigma$. First, linearize $\mathbb{P}$ about $z(j)=z^{*}$ and $I(j)=I^{*},\,r(j)=r^{*}$ as
@@
where $\mathcal{A}\in\mathbb{R}^{5\times 5}$, $\mathcal{B}\in\mathbb{R}^{5\times 2}$ may be computed numerically [khandelwal_discrete_2025]. If the pair $(\mathcal{A},\mathcal{B})$ is controllable, $\mathcal{O}^{*}$ is asymptotically stabilized by the discrete feedback
@@
Remark 6
Stabilization of a periodic orbit for initial conditions outside the domain of linearization of the desired orbit $\mathcal{O}^{*}$ may be realized by choosing a sequence of fixed points $z^{*}\equiv z^{*}(j)$ on $\Sigma$, and recomputing $\mathcal{A}$, $\mathcal{B}$, and $\mathcal{K}$ to gradually guide the system trajectory to $\mathcal{O}^{*}$ - see [khandelwal_maneuvering_2024].
@@ VI-A Enforcing the DVHC
We choose a DVHC (12) of the form
$\Phi(\theta_{k})=\begin{bmatrix}\alpha\tan\theta_{k}\\ \beta-\sigma\alpha\tan\theta_{k}\end{bmatrix},\quad\alpha=0.6131,\,\,\beta=3$ (59)
@@
$\theta^{\mathrm{odd}}=\pi/9,\,\,\theta^{\mathrm{even}}=7\pi/9\ \Rightarrow\ {\Delta\theta^{*}}=2\pi/3,\,\,\sigma=0.7779$ (60)
The choices in (59) and (60) ensure that the DZD in (38) permits stable, 2-periodic orbits - see Theorem 1. The control inputs are obtained using (27) and (28) with
@@
We consider the initial conditions
$\begin{bmatrix}q^{T}&\dot{q}^{T}\end{bmatrix}^{T}=\begin{bmatrix}0.5&2.4&0.3491&0.9&-3.2&-5.3\end{bmatrix}^{T}$ (62)
for which $\theta=\theta_{1}=\theta^{\mathrm{odd}}$ and $\rho_{1}\neq 0$. The simulation results are shown in Fig. 2 for $k=1$ through $k=20$, corresponding to a duration of approximately $10.48$ s. The components of $\rho_{k}$ are shown in Fig. 2(a)-(b), and the components of $D\rho_{k}$ are shown in Fig. 2(c)-(d). The plots demonstrate that the system trajectory converges to $\rho_{k}=0$ exponentially. As the DVHC is enforced, the system settles to a stable, 2-periodic juggling motion with
$\omega_{k}=\begin{cases}-3.1479,\ k\,\,\mathrm{odd}\\ \ \ 4.7045,\ k\,\,\mathrm{even}\end{cases}\!\!\!\!\!\!,\,\delta_{k}=\begin{cases}0.4452,\ k\,\,\mathrm{odd}\\ 0.6653,\ k\,\,\mathrm{even}\end{cases}$ (63)
which agree with (43) and (46). The values of $\omega_{k}$ and $\delta_{k}$ are plotted in Fig. 2(e) and Fig. 2(f). The control inputs $I_{k}$ and $r_{k}$ are shown in Fig. 2(g) and Fig. 2(h) respectively. For the 2-periodic juggling motion in (63), $I_{k}=0.8086$ Ns, $r_{k}=0.0202$ m for $k$ odd, and $I_{k}=-0.4302$ Ns, $r_{k}=0.0380$ m for $k$ even, in accordance with (47) and (48). The trajectory of the center-of-mass of the devil-stick corresponding to these results is shown in Fig. 3(a).
[…]
Figure 3: Trajectory of the center-of-mass of the devil-stick corresponding to the results in (a) Fig. 2 and (b) Fig. 4.
VI-B Stabilization of a Periodic Orbit
[…]
Figure 4: Orbital stabilization of 2-periodic juggling of a devil-stick from arbitrary initial conditions; $\delta_{k}=0.5442$ s $\forall k$ corresponding to stable juggling: (a)-(b) show the components of $\rho_{k}$, (c)-(d) show the components of $D\rho_{k}$, (e) shows the pre-impact angular velocity $\omega_{k}$, (f) shows the time-of-flight $\delta_{k}$ (g) shows the applied impulse $I_{k}$, and (h) shows the point of application $r_{k}$ of the impulsive force.
We now demonstrate the effectiveness of the controller described in Section V in stabilizing a desired orbit. We consider the DVHC defined by (59) and (60). Of the infinitely many stable, 2-periodic orbits, we stabilize the orbit for which $\omega^{*}=-3.8483$ rad/s, obtained from (49). This results in $\delta_{k}=0.5442$ s $\forall k$ corresponding to stable juggling. This orbit can be described using (50) as
$\begin{split}\mathcal{O}^{*}=\{(q,\dot{q}):h(k)&=\Phi(\theta_{k}),v(k)=\Psi(\theta_{k},\omega_{k}),\\ \omega_{k}&=-3.8483\,\,\forall\,\theta_{k}=\pi/9\}\end{split}$ (64)
$\Sigma=\{(q,\dot{q})\in\mathbb{R}^{2}\times S^{1}\times\mathbb{R}^{3}:\theta=\pi/9,\omega<0\}$ (65)
$z^{*}$ $=\begin{bmatrix}0.2232&2.8264&1.3553&-3.7238&-3.8483\end{bmatrix}^{T}$
$I^{*}$ $=0.7926,\quad r^{*}=0.0202$ (66)
The matrices $\mathcal{A}$ and $\mathcal{B}$ in (55) are obtained numerically as
$\begin{split}\mathcal{A}&=\begin{bmatrix}0.2500&-0.0000&-0.0000&-0.0000&-0.0000\\ 0.0000&0.2500&0.0000&-0.0000&0.0000\\ 0.0422&0.3498&0.6976&0.2539&-0.0000\\ 0.5976&-0.0428&0.8307&0.3025&-0.0000\\ -1.4247&-0.9931&-1.9809&-0.7209&0.0000\end{bmatrix}\\ \mathcal{B}&=\begin{bmatrix}0.0000&3.6387&-8.4460&-16.8780&23.9812\\ 30.4410&26.3808&-147.1159&-193.4499&282.1914\end{bmatrix}^{T}\end{split}$
$\mathcal{K}=\begin{bmatrix}0.1048&-0.0277&0.0267&0.0097&-0.0000\\ -0.0048&0.0046&0.0037&0.0014&0.0000\end{bmatrix}$
The same initial conditions (62) are used, and the simulation results are shown in Fig. 4 for $k=1$ through $k=20$, a duration of approximately $10.49$ s. The components of $\rho_{k}$ are shown in Fig. 4(a)-(b), and the components of $D\rho_{k}$ are shown in Fig. 4(c)-(d). It is seen that the system trajectory converges to $\rho_{k}=0$, and $D\rho_{k}\to 0$ as $\rho_{k}\to 0$. The values of $\omega_{k}$, shown in Fig. 4(e), converge to values in agreement with (49), $\omega^{*}=-3.8483$ rad/s, indicating stabilization of the orbit in (64). The time-of-flight, plotted in Fig. 4(f), converges to the constant value $\delta_{k}=0.5442$ s. The control inputs $I_{k}$ and $r_{k}$ are shown in Fig. 4(g) and Fig. 4(h) respectively. The inputs are obtained from (58) for $j=1,2,\dots,7$, i.e., $k=1,3,\dots,13$. The controller $u(j)$ is inactive for $j>7$ as the system trajectory is sufficiently close to $\mathcal{O}^{*}$. The inputs converge to $I_{k}=0.7926$ Ns, $r_{k}=0.0202$ m for $k$ odd, and $I_{k}=-0.4217$ Ns, $r_{k}=0.0380$ m for $k$ even in accordance with (47) and (48). The trajectory of the center-of-mass of the devil-stick corresponding to these results is shown in Fig. 3(b).
VII Conclusion
The problem of juggling a devil-stick between two angular configurations was addressed using DVHCs. This paper extends the work in [khandelwal_discrete_2025], which considered rotations of the passive coordinate to realize propeller motion, to oscillations of the passive coordinate to realize planar juggling. The DVHC approach results in a rich set of juggling motions that are not necessarily symmetric about the vertical axis; the results for planar symmetric juggling in [kant_non-prehensile_2021] are a special case. It is shown that DVHCs are an effective tool for design and control of the trajectory of an object juggled using impacts. Similar to the zero dynamics induced by VHCs, the DZD can be used to infer the nonlinear stability characteristics of the hybrid system subject to DVHCs, without relying on linearization about the fixed point of a Poincaré map [kant_juggling_2022]. Future work will focus on control design for the robot manipulator considering unilateral constraints and impact laws [zavala-rio_control_1999, brogliato_control_2000], robustness analysis of the system, and extension of the DVHC approach to general underactuated systems controlled using impulsive inputs.
Juggling between Symmetric Orientations
-A Enforcing the DVHC
For the same physical parameters of the devil-stick, we present simulation results for the special case when $\theta^{\mathrm{even}}=\pi-\theta^{\mathrm{odd}}$, i.e., the devil-stick is juggled between orientations symmetric about the vertical. We consider the same DVHC as in (59), with the values of $\theta^{\mathrm{odd}}$ and $\theta^{\mathrm{even}}$ chosen to be
$\theta^{\mathrm{odd}}=\pi/6,\,\,\theta^{\mathrm{even}}=5\pi/6\ \Rightarrow\ {\Delta\theta^{*}}=2\pi/3,\,\,\sigma=0$ (67)
Again, the choices in (59) and (67) ensure that the DZD in (38) permits stable, 2-periodic orbits - see Theorem 1 and Corollary 1. The control inputs are obtained using (27) and (28) with $\lambda$ in (61).
[…]
Figure 5: Stabilization of 2-periodic juggling of a devil-stick between orientations symmetric about the vertical from arbitrary initial conditions: (a)-(b) show the components of $\rho_{k}$, (c)-(d) show the components of $D\rho_{k}$, (e) shows the pre-impact angular velocity $\omega_{k}$, (f) shows the time-of-flight $\delta_{k}$ (g) shows the applied impulse $I_{k}$, and (h) shows the point of application $r_{k}$ of the impulsive force.
@@ VII Conclusion
which agree with (43) and (46). The values of $\omega_{k}$ and $\delta_{k}$ are plotted in Fig. 5(e) and Fig. 5(f). The control inputs $I_{k}$ and $r_{k}$ are shown in Fig. 5(g) and Fig. 5(h) respectively. For the 2-periodic juggling motion in (69), $I_{k}=-(-1)^{k}\times 0.5890$ Ns and $r_{k}=0.0308$ m, in accordance with (47) and (48) - see Remark 7.
Remark 7
For $\theta^{\mathrm{even}}=\pi-\theta^{\mathrm{odd}}$, (47) indicates that the magnitude of impulse applied at either symmetric orientation of the devil-stick is the same, with opposite signs, and (48) implies that the location of application of the impulse is identical for either symmetric orientation of the devil-stick.
The trajectory of the center-of-mass of the devil-stick corresponding to these results is shown in Fig. 6(a).
@@
Figure 6: Trajectory of the center-of-mass of the devil-stick corresponding to the results in (a) Fig. 5 and (b) Fig. 7 showing planar symmetric juggling.
-B Stabilization of a Periodic Orbit
@@
We now demonstrate the effectiveness of the controller described in Section V in stabilizing a desired orbit, specifically the orbit corresponding to planar symmetric juggling in [kant_non-prehensile_2021, kant_juggling_2022]. This juggling motion is symmetric in all states about the vertical axis. We consider the DVHC defined by (59) and (67). Of the infinitely many stable, 2-periodic orbits, we choose to stabilize the orbit for which $\omega^{*}=-4.1888$ rad/s, obtained from (49). This results in $\delta_{k}=0.5$ s $\forall k$ corresponding to stable juggling. This orbit can be described using (50) as
$\begin{split}\mathcal{O}^{*}=\{(q,\dot{q}):h(k)&=\Phi(\theta_{k}),v(k)=\Psi(\theta_{k},\omega_{k}),\\ \omega_{k}&=-4.1888\,\,\forall\,\theta_{k}=\pi/6\}\end{split}$ (70)
@@
$\Sigma=\{(q,\dot{q})\in\mathbb{R}^{2}\times S^{1}\times\mathbb{R}^{3}:\theta=\pi/6,\omega<0\}$ (71)
@@
The matrices $\mathcal{A}$ and $\mathcal{B}$ in (55) are obtained numerically as
@@
The same initial conditions (68) are used, and the simulation results are shown in Fig. 7 for $k=1$ through $k=20$, a duration of approximately $9.59$ s. The components of $\rho_{k}$ are shown in Fig. 7(a)-(b), and the components of $D\rho_{k}$ are shown in Fig. 7(c)-(d). It is seen that the system trajectory converges to $\rho_{k}=0$, and $D\rho_{k}\to 0$ as $\rho_{k}\to 0$. The values of $\omega_{k}$, shown in Fig. 7(e), converge to values in agreement with (49), $\omega^{*}=-4.1888$ rad/s, indicating stabilization of the orbit in (70). The time-of-flight, plotted in Fig. 7(f), converges to the constant value $\delta_{k}=0.5$ s. The control inputs $I_{k}$ and $r_{k}$ are shown in Fig. 7(g) and Fig. 7(h) respectively. The inputs are obtained from (58) for $j=1,2,\dots,7$, i.e., $k=1,3,\dots,13$. The controller $u(j)$ is inactive for $j>7$ as the system trajectory is sufficiently close to $\mathcal{O}^{*}$. For planar symmetric juggling, $I_{k}=-(-1)^{k}\times 0.5664$ Ns and $r_{k}=0.0308$ m in accordance with (47) and (48) - see Remark 7. The trajectory of the center-of-mass of the devil-stick corresponding to these results is shown in Fig. 6(b).